OpenStudy (anonymous):

Find a_{4} for the arithmetic series. S_{4}=-36 and a_{1}=9

ganeshie8 (ganeshie8):

use sum formula : \(\large S_n = \frac{n}{2}(a_1 + a_n)\)

ganeshie8 (ganeshie8):

\(\large S_4 = \frac{4}{2}(a_1 + a_4)\)

ganeshie8 (ganeshie8):

plugin known given values

ganeshie8 (ganeshie8):

\(\large -36 = \frac{4}{2}(9 + a_4)\)

ganeshie8 (ganeshie8):

solve \(a_4\)